# Representative Action Selection for Large Action Space Meta-Bandits

### v3

@@ Appendix F Proof of Theorem 9
Finally, note that equation 15 is a relaxation of equation 14. While the optimum of the original problem is upper bounded by that of the relaxed problem, the optimal solution to the relaxed problem also lies within the feasible region of the original problem. Therefore, the maximum of equation 14 is also attained when $q$ is uniform, with the maximum value being $\left(1-\frac{1}{|\mathcal{R}|}\right)^{K}$. ∎
Statement of Theorem 9
Under Assumption 3, there exists a point $a_{0}$ and a constant $M>0$ such that $\mathcal{A}_{\operatorname*{full}}\subset B(a_{0},M)$, a closed Euclidean ball of radius $M$ centered at $a_{0}$. Let the action space have dimension $n\in\mathbb{N}$, and fix a constant $0<\epsilon<M$. Let $\mathcal{A}$ be the output of Algorithm 1. For the same constant $C>0$ in Theorem 6, we have:
$\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\leq 2\epsilon\sqrt{n}+C\epsilon\sqrt{\log N(\mathcal{A}_{\operatorname*{full}},\epsilon)},\quad\text{where }K\geq\frac{1}{2}\left(\frac{M^{2}N(\mathcal{A}_{\operatorname*{full}},\epsilon)}{\epsilon^{2}e}-1\right).$
@@
By Lemma 15, we can shift the action space to be centered at the origin. So, without loss of generality, we assume that $\mathcal{A}_{\operatorname*{full}}\subset M\cdot B_{2}^{n}$, the scaled unit Euclidean ball in $\mathbb{R}^{n}$. Then,
@@
where the first inequality follows from $r_{\ell}\subset B(a_{\ell},\epsilon)$ by construction; the equality uses Lemma 15 to shift; and the final bound uses Claim 3 from Supplementary I. Then,
@@
$\leq$ $\epsilon\sqrt{n}+C\epsilon\sqrt{\log m}+\sqrt{\frac{m}{2K+1}}\left(\frac{2K}{2K+1}\right)^{K}\cdot M\sqrt{n}$
$\leq$ $\epsilon\sqrt{n}+C\epsilon\sqrt{\log m}+\epsilon\sqrt{n},$
where the first inequality follows from the algorithm-dependent upper bound in Theorem 7. The second inequality follows from Equations equation 16 and equation 17, and the term $\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]$ is bounded by $\frac{m}{2K+1}\left(\frac{2K}{2K+1}\right)^{2K}$ (see Lemma 17). The last inequality comes from
@@
Finally, since $\mathcal{A}_{\operatorname*{full}}\subset MB_{2}^{n}$, the covering number satisfies $N(\mathcal{A}_{\operatorname*{full}},\epsilon)\leq C^{\prime}(M/\epsilon)^{n}$ for some constant $C^{\prime}>0$; see Proposition 4.2.12 of [37]. Since $\epsilon\sqrt{\log(M/\epsilon)}\to 0$ as $\epsilon\to 0^{+}$, it follows that $\epsilon\sqrt{\log N(\mathcal{A}_{\operatorname*{full}},\epsilon)}\to 0$ as $\epsilon\to 0^{+}$. ∎
Appendix G Proof of Theorem 10
To derive a matching lower bound, we decompose the regret of Algorithm 1 into that from a reference subset and from missing clusters, using Lemma 16:
Statement of Theorem 10
Consider a partition $\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, where each cluster contains more than one action and satisfies Assumption 5. Let $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$. Let $\mathcal{A}$ be the output of Algorithm 1. Then, for the same constant $C$ in Theorem 6 and another constant $c>0$, for any reference subset $\mathcal{A}^{\prime}$:
$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\geq&c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\mathbb{E}_{\theta}\left[\operatorname*{Regret}\textnormal{of }\mathcal{A}^{\prime}\right]\\ \geq&c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\left(\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log|\mathcal{R}|}\right).\end{split}$
@@ Appendix G Proof of Theorem 10
Choose an arbitrary point $a_{\ell}\in r_{\ell}$ as the representative, such that the subset $\mathcal{A}^{\prime}=\{a_{\ell}\}_{\ell\leq m}$ forms a reference subset of Definition 1. Fix $\ell$. As in Lemma 12, define a Gaussian process $(Z_{a})_{a\in r_{\ell}}$ by setting $Z_{a}:=\mu_{a}-\mu_{a_{\ell}}$, and define random variable $Y_{\ell}:=\sup_{a\in r_{\ell}}Z_{a}$. The key idea is that $\operatorname*{Regret}\geq Y_{\ell}$ whenever $r_{\ell}\cap\mathcal{A}=\emptyset$ and $a^{*}(\theta)\in r_{\ell}$, such that
@@
where the inequality follows from Assumption 5. The last equality follows from $Y_{\ell}$ is independent from $r_{\ell}\cap\mathcal{A}\!=\!\emptyset$. Further,
$\begin{split}\mathbb{E}\left[\operatorname*{Regret}\right]=&\sum_{r\in\mathcal{R}}\Pr[r\cap\mathcal{A}\neq\emptyset,a^{*}(\theta)\in r]\cdot\mathbb{E}\left[\operatorname*{Regret}\Big|r\cap\mathcal{A}\neq\emptyset,a^{*}(\theta)\in r\right]\\ &+\sum_{r\in\mathcal{R}}\Pr[r\cap\mathcal{A}=\emptyset,a^{*}(\theta)\in r]\cdot\mathbb{E}\left[\operatorname*{Regret}\Big|r\cap\mathcal{A}=\emptyset,a^{*}(\theta)\in r\right]\\ \geq&\sum_{r\in\mathcal{R}}q(r)\cdot(1-q(r))^{K}\cdot\mathbb{E}\left[\operatorname*{Regret}\Big|r\cap\mathcal{A}=\emptyset,a^{*}(\theta)\in r\right],\\ \geq&\sum_{\ell\leq m}q(r_{\ell})\cdot(1-q(r_{\ell}))^{K}\cdot\mathbb{E}\left[Y_{\ell}\Big|a^{*}(\theta)\in r_{\ell}\right]\\ \geq&c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\left(\sum_{\ell\leq m}q(r_{\ell})\cdot\mathbb{E}^{2}\left[Y_{\ell}\Big|a^{*}(\theta)\in r_{\ell}\right]\right)^{1/2}\\ \geq&c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\mathbb{E}_{\theta}\left[\operatorname*{Regret}\textnormal{of }\mathcal{A}^{\prime}\right]\\ \geq&c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\left(\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log|\mathcal{R}|}\right)\end{split}$
where the equality uses the tower rule. The first inequality holds because the regret is lower bounded by zero when $r\cap\mathcal{A}\neq\emptyset$ and $a^{*}(\theta)\in r$, and Lemma 16. The second inequality follows from equation 18. The third inequality uses the Pólya-Szegő inequality [11], corresponding to a constant $c>0$:
@@
Note that $m_{2}$ is positive due to the assumption that each cluster contains more than one action. The fourth inequality follows from Jensen’s inequality that for a random variable $X$, $\mathbb{E}[X^{2}]\geq\mathbb{E}[X]^{2}$, and Lemma 12. Since the reference subset $\mathcal{A}^{\prime}$ is arbitrary, the lower bound $c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\mathbb{E}_{\theta}\left[\operatorname*{Regret}\textnormal{of }\mathcal{A}^{\prime}\right]$ holds for any reference subset. The last inequality uses Theorem 6. ∎
Appendix H Proof of Theorem 11
Definition 2.
@@ Appendix H Proof of Theorem 11
Let $\gamma_{2}(\mathcal{S})$ be the Talagrand’s chaining functional [34] for a set $\mathcal{S}\in\mathbb{R}^{n}$, $n\in\mathbb{N}\cup\{+\infty\}$ and Euclidean norm. It is well-known that under the assumption that $\{\mu_{a}\}_{a\in\mathcal{S}}$ is a Gaussian process, this gives the tightest bound, for a universal constant $L$:
@@
The second inequality uses the fact that $\{\mu_{a}\}_{a\in r_{\ell}}$ is a mean-zero sub-Gaussian process and applies Lemma 18. The last inequality uses the definition of $\epsilon$, and $c>0$ is a constant.
@@
Further, using Lemma 13, we have for some absolute constant $C>0$:
@@
$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq&C\sqrt{\log m}\cdot\max_{\ell\leq m}\gamma_{2}(r_{\ell})\\ &+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2}.\end{split}$
@@
If $r_{\ell}\cap\mathcal{A}\neq\emptyset$, define $a_{\ell}\in r_{\ell}\cap\mathcal{A}$. If $r_{\ell}\cap\mathcal{A}\!=\!\emptyset$, choose an arbitrary point $a_{\ell}\in r_{\ell}$ as the representative. The set $\mathcal{A}^{\prime}:=\{a_{\ell}\}_{\ell\leq m}$ forms a reference subset of Definition 1.
@@
Following the same reasoning used in the proof of Theorem 7:
@@
The first inequality follows from equation 13 in the proof of Theorem 7. The last inequality uses Lemma 19. ∎
Appendix I Properties of Gaussian width
@@
Claim 1: $\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]=\mathbb{E}\left[\max_{a^{\prime}\in-\mathcal{S}}\langle a^{\prime},\theta\rangle\right]$.
@@
Claim 2: $\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]\leq\frac{1}{2}\mathbb{E}\left[\max_{a,a^{\prime}\in\mathcal{S}}\langle a-a^{\prime},\theta\rangle\right]$.
@@
where the first equality uses Claim 1, the second equality uses the definition of $a^{*}(-\mathcal{S},\theta)$. The third equality uses linearity of expectation. The inequality uses that the vector $a^{*}(\mathcal{S},\theta)+a^{*}(-\mathcal{S},\theta)$ belongs to the set of vectors $\{a-a^{\prime}:a,a^{\prime}\in\mathcal{S}\}$. In fact, one can prove the equality that $\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]=\frac{1}{2}\mathbb{E}\left[\max_{a,a^{\prime}\in\mathcal{S}}\langle a-a^{\prime},\theta\rangle\right]$, but we only need inequality to prove the next claim.
Claim 3: $\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]\leq\frac{\operatorname{diam}(\mathcal{S})}{2}\cdot\sqrt{n}$.
@@
where the first equality uses the definition of $\mu_{a}$. The first inequality uses Claim 2. The second inequality uses Cauchy-Schwarz inequality. The third inequality uses the definition of $\operatorname{diam}(\cdot)$. The last inequality uses $\mathbb{E}\|\theta\|_{2}\leq\sqrt{n}$.
Appendix J Bounds of expected maximum of Gaussian
@@
Let $X_{1},\dots,X_{N}$ be i.i.d. $\mathcal{N}(0,\sigma^{2})$ random variables, then according to [17]:
@@
Appendix K The effect of clustering structure on regret
@@
Appendix L Varying-dependence actions with RBF/Gibbs kernels
We study the effect of varying action dependence using a Gaussian process with a kernel. To control the degree of dependence, we use stationary RBF kernel and non-stationary Gibbs kernel [39].

### v4

@@ Appendix C Extra Details of Experiments
Finally, note that (21) is a relaxation of (20). While the optimum of the original problem is upper bounded by that of the relaxed problem, the optimal solution to the relaxed problem also lies within the feasible region of the original problem. Therefore, the maximum of (20) is also attained when $q$ is uniform, with the maximum value being $\left(1-\frac{1}{m}\right)^{K}$. ∎
Assumption G.2.
The support of measure $q$ is compact.
Statement of Theorem 4.9
Under Assumption G.2, there exists a point $a_{0}$ and a constant $M>0$ such that $\mathcal{A}_{\operatorname*{full}}\subset B(a_{0},M)$, a closed Euclidean ball of radius $M$ centered at $a_{0}$. Let the action space have dimension $n\in\mathbb{N}$, and fix a constant $0<\epsilon<M$. Let $\mathcal{A}$ be the output of Algorithm 1. For the same constant $C>0$ in Theorem 4.2, we have:
$\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\leq 2\epsilon\min\{\sqrt{n},\sqrt{\frac{1}{2}\log|\mathcal{A}_{\operatorname*{full}}|}+4\}+C\epsilon\sqrt{\log N(\mathcal{A}_{\operatorname*{full}},\epsilon)},\quad\text{where }K\geq\frac{1}{2}\left(\frac{M^{2}N(\mathcal{A}_{\operatorname*{full}},\epsilon)}{\epsilon^{2}e}-1\right).$
@@
Step 1 bounding by dimension $n$:
By Lemma F.1, we can shift the action space to be centered at the origin. So, without loss of generality, we assume that $\mathcal{A}_{\operatorname*{full}}\subset M\cdot B_{2}^{n}$, the scaled unit Euclidean ball in $\mathbb{R}^{n}$. Then,
@@
where the first inequality follows from $r_{\ell}\subset B(a_{\ell},\epsilon)$ by construction; the equality uses Lemma F.1 to shift; and the final bound uses Property 3 from Supplementary J.
Step 2 bounding by cardinality $|\mathcal{A}_{\operatorname*{full}}|$:
By the definition of variance:
$\begin{split}\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}^{2}\right]=&\left(\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}\right]\right)^{2}+\text{Var}\left(\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}\right)\\ \leq&\left(\frac{\operatorname{diam}(\mathcal{A}_{\operatorname*{full}})}{2}\sqrt{2\log|\mathcal{A}_{\operatorname*{full}}|}\right)^{2}+4\operatorname{diam}(\mathcal{A}_{\operatorname*{full}})^{2}\\ =&M^{2}\left(\sqrt{\frac{1}{2}\log|\mathcal{A}_{\operatorname*{full}}|}+4\right)^{2},\end{split}$ (24)
where the inequality uses Supplementary K to bound the square of Gaussian width, and Lemma E.3 to bound the variance. The last equality uses the definition of $M$ and simple rearrangement.
Also,
$\max_{\ell\leq m}\mathbb{E}_{\theta}\left[\max_{a\in r_{\ell}}\mu_{a}\right]\leq\max_{\ell\leq m}\frac{\operatorname{diam}(r_{\ell})}{2}\sqrt{2\log|r_{\ell}|}\leq\frac{\epsilon}{2}\sqrt{2\log|\mathcal{A}_{\operatorname*{full}}|},$ (25)
where the left inequality uses Supplementary K, the right inequality uses the definition of $\epsilon$.
Then,
@@
$\leq$ $\epsilon\min\{\sqrt{n},\sqrt{\frac{1}{2}\log|\mathcal{A}_{\operatorname*{full}}|}\}+C\epsilon\sqrt{\log m}+\sqrt{\frac{m}{2K+1}}\left(\frac{2K}{2K+1}\right)^{K}M\;\min\{\sqrt{n},\sqrt{\frac{1}{2}\log|\mathcal{A}_{\operatorname*{full}}|}+4\}$
$\leq$ $\epsilon\min\{\sqrt{n},\sqrt{\frac{1}{2}\log|\mathcal{A}_{\operatorname*{full}}|}\}+C\epsilon\sqrt{\log m}+\epsilon\min\{\sqrt{n},\sqrt{\frac{1}{2}\log|\mathcal{A}_{\operatorname*{full}}|}+4\}$
$\leq$ $2\epsilon\min\{\sqrt{n},\sqrt{\frac{1}{2}\log|\mathcal{A}_{\operatorname*{full}}|}+4\}+C\epsilon\sqrt{\log N(\mathcal{A}_{\operatorname*{full}},\epsilon)},$
where the first inequality follows from the algorithm-dependent upper bound in Theorem 4.5. The second inequality follows from Equations (22)-(25), and the term $\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]$ is bounded by $\frac{m}{2K+1}\left(\frac{2K}{2K+1}\right)^{2K}$ (see Lemma G.1). The third inequality comes from
@@
Finally, since $\mathcal{A}_{\operatorname*{full}}\subset MB_{2}^{n}$, the covering number satisfies $N(\mathcal{A}_{\operatorname*{full}},\epsilon)\leq C^{\prime}(M/\epsilon)^{n}$ for some constant $C^{\prime}>0$; see Proposition 4.2.12 of (Vershynin, 2018). Since $\epsilon\sqrt{\log(M/\epsilon)}\to 0$ as $\epsilon\to 0^{+}$, it follows that $\epsilon\sqrt{\log N(\mathcal{A}_{\operatorname*{full}},\epsilon)}\to 0$ as $\epsilon\to 0^{+}$. ∎
Appendix H Proof of Theorem B.1
To derive a matching lower bound, we decompose the regret of Algorithm 1 into that from a reference subset and from missing clusters, using Lemma F.2:
Statement of Theorem B.1
Consider a partition $\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, where each cluster contains more than one action and satisfies Assumption 4.3. Let $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$. Let $\mathcal{A}$ be the output of Algorithm 1. Then, for the same constant $C$ in Theorem 4.2 and another constant $c>0$, for any reference subset $\mathcal{A}^{\prime}$:
$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\geq&c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\mathbb{E}_{\theta}\left[\operatorname*{Regret}\textnormal{of }\mathcal{A}^{\prime}\right]\\ \geq&c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\left(\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log m}\right).\end{split}$
@@
Choose an arbitrary point $a_{\ell}\in r_{\ell}$ as the representative, such that the subset $\mathcal{A}^{\prime}=\{a_{\ell}\}_{\ell\leq m}$ forms a reference subset of Definition 4.1. Fix $\ell$. As in Lemma E.1, define a Gaussian process $(Z_{a})_{a\in r_{\ell}}$ by setting $Z_{a}:=\mu_{a}-\mu_{a_{\ell}}$, and define random variable $Y_{\ell}:=\sup_{a\in r_{\ell}}Z_{a}$. The key idea is that $\operatorname*{Regret}\geq Y_{\ell}$ whenever $r_{\ell}\cap\mathcal{A}=\emptyset$ and $a^{*}(\theta)\in r_{\ell}$, such that
@@
where the inequality follows from Assumption 4.3. The last equality follows from $Y_{\ell}$ is independent from $r_{\ell}\cap\mathcal{A}\!=\!\emptyset$. Further,
$\begin{split}\mathbb{E}\left[\operatorname*{Regret}\right]=&\sum_{r\in\mathcal{R}}\Pr[r\cap\mathcal{A}\neq\emptyset,a^{*}(\theta)\in r]\cdot\mathbb{E}\left[\operatorname*{Regret}\Big|r\cap\mathcal{A}\neq\emptyset,a^{*}(\theta)\in r\right]\\ &+\sum_{r\in\mathcal{R}}\Pr[r\cap\mathcal{A}=\emptyset,a^{*}(\theta)\in r]\cdot\mathbb{E}\left[\operatorname*{Regret}\Big|r\cap\mathcal{A}=\emptyset,a^{*}(\theta)\in r\right]\\ \geq&\sum_{r\in\mathcal{R}}q(r)\cdot(1-q(r))^{K}\cdot\mathbb{E}\left[\operatorname*{Regret}\Big|r\cap\mathcal{A}=\emptyset,a^{*}(\theta)\in r\right],\\ \geq&\sum_{\ell\leq m}q(r_{\ell})\cdot(1-q(r_{\ell}))^{K}\cdot\mathbb{E}\left[Y_{\ell}\Big|a^{*}(\theta)\in r_{\ell}\right]\\ \geq&c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\left(\sum_{\ell\leq m}q(r_{\ell})\cdot\mathbb{E}^{2}\left[Y_{\ell}\Big|a^{*}(\theta)\in r_{\ell}\right]\right)^{1/2}\\ \geq&c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\mathbb{E}_{\theta}\left[\operatorname*{Regret}\textnormal{of }\mathcal{A}^{\prime}\right]\\ \geq&c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\left(\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log m}\right)\end{split}$
where the equality uses the tower rule. The first inequality holds because the regret is lower bounded by zero when $r\cap\mathcal{A}\neq\emptyset$ and $a^{*}(\theta)\in r$, and Lemma F.2. The second inequality follows from (26). The third inequality uses the Pólya-Szegő inequality (Dragomir, 2004), corresponding to a constant $c>0$:
@@
Note that $m_{2}$ is positive due to the assumption that each cluster contains more than one action. The fourth inequality follows from Jensen’s inequality that for a random variable $X$, $\mathbb{E}[X^{2}]\geq\mathbb{E}[X]^{2}$, and Lemma E.1. Since the reference subset $\mathcal{A}^{\prime}$ is arbitrary, the lower bound $c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\mathbb{E}_{\theta}\left[\operatorname*{Regret}\textnormal{of }\mathcal{A}^{\prime}\right]$ holds for any reference subset. The last inequality uses Theorem 4.2. ∎
Appendix I Regret Bound for Sub-Gaussian Processes
The Gaussian process assumption can be relaxed to a sub-Gaussian one, where the linear structure in Equation (5) is not required, and the random process $\{\mu_{a}\}_{a\in\mathcal{S}}$ satisfies the following increment condition:
$\forall u>0,\;\Pr\left[|\mu_{a}\!-\!\mu_{a^{\prime}}|\geq u\right]\leq 2\exp\left(-\frac{u^{2}}{2\|a\!-\!a^{\prime}\|_{2}^{2}}\right).$ (27)
Let $\gamma_{2}(\mathcal{S})$ denote Talagrand’s $\gamma_{2}$-functional (Talagrand, 2014) for a set $\mathcal{S}\subset\mathbb{R}^{n}$, $n\in\mathbb{N}$, equipped with the Euclidean norm. It is known that for any sub-Gaussian process indexed by $\mathcal{S}$, there exists a universal constant $c>0$ such that $\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]\leq c\,\gamma_{2}(\mathcal{S}).$
Using this functional, we establish the following regret bound for our algorithm:
Theorem I.1.
Let $\{\mu_{a}\}_{a\in\mathcal{A}_{\mathrm{full}}}$ be a mean-zero sub-Gaussian process satisfying Equation (27). Let $\mathcal{A}$ be the output of Algorithm 1. Then, for some constant $C>0$,
$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\mathrm{Regret}]\leq\;C\sqrt{\log m}\cdot\max_{\ell\leq m}\gamma_{2}(r_{\ell})+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\mathrm{full}}}\mu^{2}_{a}\right]\right)^{1/2}.\end{split}$
Definition I.2.
@@ Appendix I Regret Bound for Sub-Gaussian Processes
Let $\gamma_{2}(\mathcal{S})$ be the Talagrand’s chaining functional (Talagrand, 2014) for a set $\mathcal{S}\in\mathbb{R}^{n}$, $n\in\mathbb{N}\cup\{+\infty\}$ and Euclidean norm. It is well-known that under the assumption that $\{\mu_{a}\}_{a\in\mathcal{S}}$ is a Gaussian process, this gives the tightest bound, for a universal constant $L$:
@@
The second inequality uses the fact that $\{\mu_{a}\}_{a\in r_{\ell}}$ is a mean-zero sub-Gaussian process and applies Lemma I.3. The last inequality uses the definition of $\epsilon$, and $c>0$ is a constant.
@@
Further, using Lemma E.2, we have for some absolute constant $C>0$:
@@
$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq&C\sqrt{\log m}\cdot\max_{\ell\leq m}\gamma_{2}(r_{\ell})+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2}.\end{split}$
@@
If $r_{\ell}\cap\mathcal{A}\neq\emptyset$, define $a_{\ell}\in r_{\ell}\cap\mathcal{A}$. If $r_{\ell}\cap\mathcal{A}\!=\!\emptyset$, choose an arbitrary point $a_{\ell}\in r_{\ell}$ as the representative. The set $\mathcal{A}^{\prime}:=\{a_{\ell}\}_{\ell\leq m}$ forms a reference subset of Definition 4.1.
@@
Following the same reasoning used in the proof of Theorem 4.5:
@@
The first inequality follows from (19) in the proof of Theorem 4.5. The last inequality uses Lemma I.4. ∎
Appendix J Properties of Gaussian width
@@
Property 1: $\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]=\mathbb{E}\left[\max_{a^{\prime}\in-\mathcal{S}}\langle a^{\prime},\theta\rangle\right]$.
@@
Property 2: $\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]\leq\frac{1}{2}\mathbb{E}\left[\max_{a,a^{\prime}\in\mathcal{S}}\langle a-a^{\prime},\theta\rangle\right]$.
@@
where the first equality uses Property 1, the second equality uses the definition of $a^{*}(-\mathcal{S},\theta)$. The third equality uses linearity of expectation. The inequality uses that the vector $a^{*}(\mathcal{S},\theta)+a^{*}(-\mathcal{S},\theta)$ belongs to the set of vectors $\{a-a^{\prime}:a,a^{\prime}\in\mathcal{S}\}$. In fact, one can prove the equality that $\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]=\frac{1}{2}\mathbb{E}\left[\max_{a,a^{\prime}\in\mathcal{S}}\langle a-a^{\prime},\theta\rangle\right]$, but we only need inequality to prove the next claim.
Property 3: $\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]\leq\frac{\operatorname{diam}(\mathcal{S})}{2}\cdot\sqrt{n}$.
@@
where the first equality uses the definition of $\mu_{a}$. The first inequality uses Property 2. The second inequality uses Cauchy-Schwarz inequality. The third inequality uses the definition of $\operatorname{diam}(\cdot)$. The last inequality uses $\mathbb{E}\|\theta\|_{2}\leq\sqrt{n}$.
Appendix K Bounds of expected maximum of Gaussian
@@
Let $X_{1},\dots,X_{N}$ be i.i.d. $\mathcal{N}(0,\sigma^{2})$ random variables, then according to (Kamath, 2015):
@@
Appendix L The effect of clustering structure on regret
@@
Appendix M Varying-dependence actions with RBF/Gibbs kernels
We study the effect of varying action dependence using a Gaussian process with a kernel. To control the degree of dependence, we use stationary RBF kernel and non-stationary Gibbs kernel (Williams & Rasmussen, 2006).